OpenStudy (anonymous):

lim ---- (x^2+4x-3) x->2

OpenStudy (anonymous):

Simply plug in 2. 2^2+4(2)-3=4+8-3=9

OpenStudy (anonymous):

so it is 9?

OpenStudy (anonymous):

Yup :D